OpenStudy (anonymous):

use substitution x+3y=-3 -2x-4y=2

OpenStudy (anonymous):

x = -3y - 3 then 2(-3y - 3 ) - 4y = 2 then 6y - 6 - 4y = 2 then 6y -4y = 2- 6 then simplify 2y = -2 divide both side by 2, 2y/2 = -2/2 and Y = -1 subsitute to x = -3y -3 equals to x = -3 (-1) - 3 And the X = 0 Y = -1 X = 0 Thats it? .. Guess?